A car moving with a speed of 50km/hr can be stopped by breaks after all least 6m. if thesame car is moving at a speed of 100km/hr the minimum stopping distance is --------?


1
Expert's answer
2021-03-29T08:59:03-0400

Let's first find the deceleration of the car:


"v^2=v_0^2+2as,""0=v_0^2+2as,""a=\dfrac{-v_0^2}{2s}=\dfrac{-(13.89\ \dfrac{m}{s})^2}{2\cdot6\ m}=-16.1\ \dfrac{m}{s^2}."

Finally, we can find the minimum stopping distance if the same car is moving at a speed of 100km/hr:


"s=\dfrac{-v_0^2}{2a}=\dfrac{-(27.78\ \dfrac{m}{s})^2}{2\cdot(-16.1\ \dfrac{m}{s^2})}=24\ m."